A bond with 10 years to maturity, has a semiannual coupon payment of $300, and a current price of $7,231.62, If the yield to maturity is 6%, what is the price of the bond in percentage of par?
Answer #1

Value of BOnd = 1 1 (1+r) Coupon * + MaturityValue (1 + r) 7

r = 0.06 / 2 = 0.03

n = 10 yearas * 2 = 20

7231.62 = \small 300 *\frac{1-\frac{1}{(1+0.03)^{20}}}{0.03} + \frac{Maturity Value}{(1+0.03)^{20}}

7231.62 = 4463.2424589 + Maturity Value * 0.55367575411

Maturity Value = (7231.62 - 4463.2424589) / 0.55367575411

Maturity Value = 5000

Value of Bond in % = 7231.62 / 5000 = 144.63%

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.